OpenStudy (anonymous):

please someone who really knows what they are doing in excel please help me. Please if you dont know then dont respond.

OpenStudy (anonymous):

all I am suppose to do for the rest of this assignment is the formula that adds the total from Jan through Jun with the monthly figures from Jul through Dec. Copy the formula down the column. Change the average formula to average the 12 months. Clear the formula from the cells that reference blanks cells. Clear the conditional formatting from all cells in Jul through Dec sales sheet. Check formulas by entering (and removing) some sample data. Change the label in H4 to Annual Total and I4 to Annual AVG. Best fit both columns.

OpenStudy (anonymous):

i know exel pretty well i will try to help could you mabey type what you need in a list form so i understand a little bit more than i can work it out

OpenStudy (anonymous):

like 1. 2. 3. 4. that kinda thing like a check list

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know i can create what you need but explaining how i did it ill try my best

OpenStudy (anonymous):

1.In H6 enter a formula that adds the total from Jan through Jun with the monthly figures from Jul through Dec.

OpenStudy (anonymous):

okay is this formula somthing in exel or somthing like 2+2

OpenStudy (ash2326):

I'm also doing it, wait for some time

OpenStudy (anonymous):

in excel

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i attached the worksheet

OpenStudy (anonymous):

when i add the formula in does it go on witch spread sheet

OpenStudy (anonymous):

im sorry if theres any frustration im just trying to figure out a lil more as i do it

OpenStudy (anonymous):

july-dec in H6

OpenStudy (anonymous):

then I have to copy the formula down the column

OpenStudy (anonymous):

do you want that formula to add on too number that you will place there for those monthly avg.?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay i understand :)

OpenStudy (anonymous):

ok these are the steps to finish the worksheet i have attached: 1. In cell H6, enter a formula that adds the total from Jan through Jun with the monthly figures from Jul through Dec. 2. Copy the formula down the column. 3. Change the average formula to average the 12 months. 4. Clear the formula from the cells that reference blanks cells. 5. Clear the conditional formatting from all cells in Jul through Dec sales sheet. 6.Check formulas by entering (and removing) some sample data. 7.Change the label in H4 to Annual Total and I4 to Annual AVG. 8.Best fit both columns.

OpenStudy (ash2326):

Tina check this one

OpenStudy (ash2326):

with sample data

OpenStudy (ash2326):

did you check it??

OpenStudy (anonymous):

it looks good but how did you do it

OpenStudy (anonymous):

however that one doesnt add to the total

OpenStudy (anonymous):

type numbers into the blanks it adds to it

OpenStudy (ash2326):

let's say for annual total, we want to sum total of Jan to jun total + july-dec in the excel sheet there was already the sum of july to dec by the formula sum(B6:G6) we need to add jan to jun total with it, so write the formula in H6 as ''=sum(B6:G6)+ ____'' after writing + sign , go to the fist sheet and click H6 there you'll get a formula as =SUM(B6:G6)+'January-June Sales'!H6 when you copy this formula for the lower cells,it'll automatically change to the respective cell numbers. to find average just put in I6"=H6/12" then copy this in the whole column.

OpenStudy (anonymous):

this one will add to the jan- july total

OpenStudy (ash2326):

did you get it tina??

OpenStudy (ash2326):

sorry, it took me so long:(

OpenStudy (anonymous):

i can not open that file for some reason

OpenStudy (ash2326):

which file

OpenStudy (ash2326):

kamuela's file is not opening

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please ash write the the formulas for me

OpenStudy (anonymous):

kamuela your file can not be opened

OpenStudy (anonymous):

please ash write the the formulas for me here

OpenStudy (anonymous):

im exhausted and going to bed its 4:30am and i still have a whole other worksheet to do plus a quiz all due sunday by 11:59p

OpenStudy (ash2326):

Total"=SUM(B6:G6)+'January-June Sales'!H6" Average"=H6/12"

OpenStudy (anonymous):

perfect ty

OpenStudy (ash2326):

type the total under total Annual and then copy in all the cells type the average under annual average and then copy in all cells

OpenStudy (ash2326):

now you sleep kid